Ideal Body Weight of 71 kg The ventilator tidal volume range is: Ideal Body weight of 138 lb IBW is kg The ventilator tidal volume range is: Ideal Body Weight of 55 kg The ventilator tidal volume range is: Ideal Body weight of 133 lb BW is The ventilator tidal volume range is: kg The peak inspiratory pressue is 47 cmH20 and the PEEP is 8 cmH20 with a set Vt of 750 mL. The tubing compliance factor is 3.4 mL/cmH20. a. what is the lost volume b. delivered volume: The peak inspiratory pressue is 52 cmH20 and the PEEP is 5 cmH20 with a set Vt of 700 mL. The tubing compliance factor is 3.4 mL/cmH20. a, what is the lost volume: b. delivered volume:
Answer #1

ANSWER

   Ventillator tidal volume

       * Patient on ventillator due to lung diseases set tidal volume 6ml/kg to 8 ml/kg.

              So

                    Formula = 6 - 8 ml /kg IBW.

                   * Ideal body weight = 71 kg.

                        Ventillator tidal volume range = 71 * 6 to 71 * 8

                                                                       = 426ml to 568 ml.

                 * Ideal body weight = 138 lb.

                            IBW in kg = 137 / 2.205 (divided by 2.205 for covert lb to kg.)

                                              = 62.14 kg.

                        Ventillator tidal volume range = 62.14 * 6 to 62.14 * 8

                                                                       = 372.8ml to 497.1 ml

                   * Ideal body weight = 55 kg.

                          Ventillator tidal volume range = 55 * 6 to 55 * 8

                                                                          = 330 ml to 440 ml.

                  * Ideal body weight = 133 lb.

                            lBW in kg = 133 / 2.205 (divided by 2.205 for convert lb to kg)

                                             = 60.3 kg.

                              Ventillator tidal volume range = 60 * 6 to 60 * 8

                                                                             = 360ml to 480ml.

* Formula for calculate lost volume

                       Formula = Compliance factor * Peak pressure.

* Formula for delivered volume.

                      Formula = Tidal volume - Lost volume.

                   As per formula

                       a) Lost volume.

                               Compliance factor = 3.4 ml /cmH2O.

                         Peak pressure = 45 cmH2O.

                               Lost volume = 3.4 * 45

                                                  = 153 ml.

                      b) Delivered volume.

                                Tidal volume = 750 ml.

                            Lost volume = 153 ml.

                             Delivered volume = 750 - 153

                                                           = 597ml.

                     a) Lost volume.

                                Compliance factor = 3.4 ml/cmH2O.

                                  Peak pressure = 52 cmH2O.

                              Lost volume = 3.4 * 52

                                                   = 176.8 ml.

                       b) Delivered volume.

                               Tidal volume = 700 ml.

                            Lost volume = 176.8 ml.

                            Delivered volume = 700 - 176.8

                                                                       = 523.2 ml.